# A MIND for Reasoning: Meta-learning for In-context Deduction

### v1

@@ Abstract
Large language models (LLMs) are increasingly evaluated on formal tasks, where strong reasoning abilities define the state of the art. However, their ability to generalize to out-of-distribution problems remains limited. In this paper, we investigate how LLMs can achieve a systematic understanding of deductive rules. Our focus is on the task of identifying the appropriate subset of premises within a knowledge base needed to derive a given hypothesis. To tackle this challenge, we propose Meta-learning for IN-context Deduction (MIND), a novel few-shot meta-learning fine-tuning approach. The goal of MIND is to enable models to generalize more effectively to unseen knowledge bases and to systematically apply inference rules. Our results show that MIND significantly improves generalization in small LMs ranging from 1.5B to 7B parameters. The benefits are especially pronounced in smaller models and low-data settings. Remarkably, small models fine-tuned with MIND outperform state-of-the-art LLMs, such as GPT-4o and o3-mini, on this task.
A MIND for Reasoning: Meta-learning for In-context Deduction
Leonardo Bertolazzi1, Manuel Vargas Guzmán2, Raffaella Bernardi3, Maciej Malicki2, Jakub Szymanik1, 1University of Trento, 2University of Warsaw, 3Free University of Bozen-Bolzano
@@ 1 Introduction
Reasoning refers to a broad set of abilities that are applied not only in formal domains, such as mathematics and logic, but also in goal-directed scenarios involving problem-solving and decision-making (Leighton, 2004). All types of reasoning share a common foundation: the capacity to reach an abstract understanding of the problem at hand. With the advent of increasingly capable large language models (LLMs), reasoning has become a central domain for evaluating and comparing these systems (Huang and Chang, 2023; Mondorf and Plank, 2024).
[…]
@@
Figure 1: Overview of a MIND episode. Given a set of premises (the knowledge base, $\mathcal{KB}$), a set of task demonstrations (or study examples, denoted by the <STUDY> tag), and a query hypothesis $x^{\mathrm{query}}$ (denoted by the <QUERY> tag) that is entailed from $\mathcal{KB}$, models must generate the minimal subset of premises $y^{\mathrm{query}}$ from which $x^{\mathrm{query}}$ can be derived. During each MIND episode, models can practice on hypothesis-premise pairs before processing the main query hypothesis. The examples show how we frame syllogistic inferences as a premise selection task.
Despite extensive training on mathematical, programming, and STEM-related data, LLMs continue to struggle in out-of-distribution (OOD) reasoning scenarios. Their performance often deteriorates on longer inference chains than those seen during training (Clark et al., 2021; Saparov et al., 2023), and they exhibit variability when evaluated with perturbed versions of the same problems (Mirzadeh et al., 2025; Gulati et al., 2024; Huang et al., 2025). In particular, LLMs can get distracted by irrelevant context, becoming unable to solve problems they could otherwise solve (Shi et al., 2023; Yoran et al., 2024). These challenges relate to broader debates surrounding generalization versus memorization in LLMs (Balloccu et al., 2024; Singh et al., 2024).
@@
Few-shot meta-learning approaches (Irie and Lake, 2024) have emerged as promising methods for inducing OOD generalization and rapid domain adaptation in LLMs. Specifically, this class of methods has proven effective in few-shot task generalization (Min et al., 2022; Chen et al., 2022), systematic generalization (Lake and Baroni, 2023), and mitigating catastrophic forgetting (Irie et al., 2025).
@@
In this work, we propose Meta-learning for IN-context Deduction (MIND), a new few-shot meta-learning fine-tuning approach for deductive reasoning. As illustrated in Figure 1, we evaluate the effectiveness of this approach using a logical reasoning task grounded in syllogistic logic (Smiley, 1973; Vargas Guzmán et al., 2024). Each problem presents a knowledge base of atomic logical statements. Models are tasked with identifying the minimal subset of premises that logically entail a given test hypothesis. This premise selection task captures a core aspect of deductive reasoning: determining which known facts are necessary and sufficient to justify a conclusion. We apply MIND to small LMs from the Qwen-2.5 family (Qwen Team, 2025), ranging from 1.5B to 7B parameters. Specifically, we assess the generalization capabilities induced by MIND, such as systematically performing inferences over unseen sets of premises, as well as over more complex (longer) or simpler (shorter) sets of premises than those encountered during training. [^1] Our code and data are available at: https://github.com/leobertolazzi/MIND.git
Our main contributions are as follows:
  • We introduce a new synthetic dataset based on syllogistic logic to study reasoning generalization in LLMs.
  • We show that MIND enables LMs to better generalize in OOD reasoning problems with particularly strong performance in smaller models and low-data regimes.
  • We demonstrate that small LMs fine-tuned with MIND can outperform state-of-the-art LLMs such as GPT-4o and o3-mini, on our premise selection task.
@@ 2.1 Syllogistic Logic
Figure 2: Example inference. Edges labeled “All-are” denote universal affirmatives (e.g., All cats are felines). The solid red edge is a universal negative (No animals are plants). From these “atomic facts” we infer No cats are tulips (dashed red edge). Formally, this is expressed as $\{Aa-b,\;Ac-d,\;Ebd\}\vDash Eac$ (Smiley, 1973).
@@
The language of syllogistic logic comprises a finite set of atomic terms $\{a,b,c,\ldots\}$ and four quantifier labels $A,E,I$, and $O$. Well-formed formulas consists of $Aab$ (“All $a$ are $b$”), $Eab$ (“No $a$ are $b$”), $Iab$ (“Some $a$ are $b$”), and $Oab$ (“Some $a$ are not $b$”). Finally, an A-chain, denoted as $Aa-b$ represents the single formula $Aab$ or a sequence of formulas $Aac_{1}$, $Ac_{1}c_{2}$, $\dots$, $Ac_{n-1}c_{n}$, $Ac_{n}b$ for $n\geq 1$. A knowledge base ($\mathcal{KB}$) is defined as a finite set of formulas (premises).
@@
We aim for models to identify the minimal set of premises in a knowledge base to derive a given hypothesis. Formally, we are interested in inferences $\mathcal{F}\vDash F$ such that ⊭$\mathcal{F}^{\prime}\not\vDash F$ for any proper subset $\mathcal{F}^{\prime}\subsetneq\mathcal{F}$. For example, $\{Abc,Abd\}\vDash Icd$ is minimal, while $\{Aab,Abc,Abd\}\vDash Icd$ is not because $Aab$ is not needed to infer the conclusion.
There are seven types of minimal syllogistic inferences. [^2] See the full list in Table 4 in Appendix A. To facilitate understanding, Figure 2 provides an intuitive representation of a type 6 inference. Further details about the syllogistic logic can be found in Appendix A.
@@ 2.2 Meta-learning in Autoregressive Models
Meta-learning, or “learning to learn”, is a paradigm that aims to enable machine learning models to acquire transferable knowledge across multiple tasks, allowing rapid adaptation to new tasks with minimal data. Among the numerous existing meta-learning frameworks (Hospedales et al., 2022), MIND is mainly inspired by Meta-learning Sequence Learners (MSL) (Irie and Lake, 2024).
@@
In standard supervised learning, data consists of a static dataset $\mathcal{D}_{\mathrm{train}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ where inputs $x_{i}$ are mapped to targets $y_{i}$ under a fixed distribution $p(x,y)$. By contrast, meta-learning organizes data into tasks (or episodes) $\mathcal{T}=(S^{\mathrm{supp}},S^{\mathrm{query}})$ drawn from $p(\mathcal{T})$, where $S^{\mathrm{supp}}=\{(x_{i},y_{i})\}_{i=1}^{K}$ is the support set containing task demonstrations, or study examples, and $S^{\mathrm{query}}=\{(x_{j},y_{j})\}_{j=1}^{M}$ is the query set for evaluation. We consider the simplest scenario where $|S^{\mathrm{query}}|=1$, containing a single example $(x^{\mathrm{query}},y^{\mathrm{query}})$. We adapt this episodic formulation to our task, as shown in Figure 1.
@@
The fundamental difference between the two paradigms appears in their optimization objectives. Standard supervised learning finds parameters $\theta^{*}$ that maximize the likelihood:
[…]
while meta-learning finds parameters $\theta^{*}$ that maximize the expected likelihood across tasks:
[…]
For autoregressive models, the probability $p_{\theta}(y^{\mathrm{query}}\mid x^{\mathrm{query}},S^{\mathrm{supp}})$ is computed by conditioning on the support set $S^{\mathrm{supp}}$ as part of the input context, formatted as a sequence of input-output pairs preceding the query. This approach forces the model to develop the capabilities of recognizing and applying task patterns from the support examples to generate appropriate query outputs.
@@ 3.1 Data Generation
As illustrated in Figure 1, we structured each datapoint in the three splits to begin with the token “knowledge base:”, followed by the full sequence of premises, separated by commas. This is immediately followed by the special tag <QUERY>, and then the token “hypothesis:”, which introduces the target hypothesis. Next comes the token “premises:”, followed by the specific comma-separated premises that entail the hypothesis. To increase variability, we applied ten random pseudoword assignments and three random permutations of premise order for each $\mathcal{KB}$, resulting in multiple variants per datapoint.
@@
We generated enough $\mathcal{KB}$s to obtain 1000 training, 5 validation, and 100 test examples for each inference type and length combination in the range from 0 to 19.666Note that some inference types (e.g., type 3) span the full range of lengths from 0 to 19, while others span only a subrange (e.g., type 2 spans from 1 to 10). See all type-length combinations within the generated $\mathcal{KB}$s in Figure 6 in Appendix B.4. This range was chosen to allow experiments with generalization to both unseen shorter and longer inferences. Full dataset statistics, including the number of generated $\mathcal{KB}$s per split, are reported in Appendix B.4.
@@
3.2 MIND
When applying meta-learning principles to the framework of syllogistic logic, we conceptualize the premises within a $\mathcal{KB}$ as atomic facts. The seven types of syllogism (as detailed in Table 4) are treated as arguments, constructed using these atomic facts, and the model’s task is to extract the minimal set of facts within a $\mathcal{KB}$ to produce a valid argument that proves the query hypothesis.
The type of systematic generalization MIND addresses involves applying the seven fixed syllogistic inferences to new, unseen sets of atomic facts. This is central to logical reasoning because logical rules are, by definition, formal: conclusions follow from premises based solely on the structure of the arguments, regardless of their specific content. Thus, successfully applying an inference type to a novel, unseen $\mathcal{KB}$ requires the model to recognize and instantiate the same formal structure with different premises. This generalization also includes variations in the number of atomic facts needed to instantiate an argument. Specifically, handling $A$-chains of varying lengths requires applying the learned inference patterns to longer or shorter instances of the same formal type.
@@ 3.2 MIND
To induce meta-learning of inference types, MIND uses a set of episodes where each episode $\mathcal{T}=(\mathcal{KB},S^{\mathrm{supp}},x^{\mathrm{query}},y^{\mathrm{query}})$. Here, $\mathcal{KB}$ is a knowledge base, $S^{\mathrm{supp}}$ is a set of study valid hypothesis-premises pairs, $x^{\mathrm{query}}$ is a valid query hypothesis, and $y^{\mathrm{query}}$ is the minimal set of premises entailing $x^{\mathrm{query}}$. Figure 1 shows a full MIND episode using indexed variables in place of pseudowords for improved readability. Importantly, we consider study examples with inferences of the same type as the query. The number of study examples we set, i.e. valid hypothesis–premise pairs, is three. In their textual translation, we add the special tags <STUDY> to indicate the beginning of the sequence of study examples. During MIND fine-tuning, models are trained to minimize the cross-entropy loss of the tokens in $y^{\mathrm{query}}$ given the input tokens from the context $(\mathcal{KB},S^{\mathrm{supp}},x^{\mathrm{query}})$.
@@
Similarly to Lake and Baroni (2023), we consider a baseline where models are not fine-tuned on episodes but on single input-output pairs $(x^{\mathrm{query}},y^{\mathrm{query}})$ preceded by a $\mathcal{KB}$. The baseline is fine-tuned to minimize the cross-entropy loss of the tokens in $y^{\mathrm{query}}$ given the input tokens from the context $(\mathcal{KB},x^{\mathrm{query}})$. To ensure a fair comparison between the meta-learning model and the baseline, we ensured that both models were fine-tuned on the exact same aggregate set of unique hypothesis-premises pairs. Specifically, the baseline was fine-tuned using a set $\mathcal{D}_{\text{baseline}}$ consisting of $(x^{\mathrm{query}},y^{\mathrm{query}})$ unique pairs. For the meta-learning approach, the corresponding set of all unique hypothesis-premises pairs encountered across all $N$ episodes $\mathcal{T}_{i}=(\mathcal{KB}_{i},S^{\mathrm{supp}}_{i},x^{\mathrm{query}}_{i},y^{\mathrm{query}}_{i})$ is given by $\mathcal{D}_{\text{meta}}=\bigcup_{i=1}^{N}(S^{\mathrm{supp}}_{i}\cup\{(x^{\mathrm{query}}_{i},y^{\mathrm{query}}_{i})\})$. We verified that $\mathcal{D}_{\text{baseline}}=\mathcal{D}_{\text{meta}}$. Moreover, since the meta-learning model processes more hypothesis-premises pairs within each episode (due to $S^{\mathrm{supp}}_{i}$), we counterbalanced this by training the baseline model for a proportionally larger number of epochs. [^7] Further details on the training regime and number of epochs for each approach are provided in Appendix C.2.

### v2

# Teaching Small Language Models to Learn Logic through Meta-Learning

@@ Abstract
Large language models (LLMs) are increasingly evaluated on reasoning tasks, yet their logical abilities remain contested. To address this, we study LLMs’ reasoning in a well-defined fragment of logic: syllogistic reasoning. We cast the problem as premise selection and construct controlled datasets to isolate logical competence. Beyond evaluation, an open challenge is enabling LLMs to acquire abstract inference patterns that generalize to novel structures. We propose to apply few-shot meta-learning to this domain, thereby encouraging models to extract rules across tasks rather than memorize patterns within tasks. Although meta-learning has been little explored in the context of logic learnability, our experiments show that it is effective: small models (1.5B–7B) fine-tuned with meta-learning demonstrate strong gains in generalization, with especially pronounced benefits in low-data regimes. These meta-learned models outperform GPT-4o and o3-mini on our syllogistic reasoning task.
Teaching Small Language Models to Learn Logic through Meta-Learning
Leonardo Bertolazzi1, Manuel Vargas Guzmán2, Raffaella Bernardi3, Maciej Malicki2, Jakub Szymanik1, 1University of Trento, 2University of Warsaw, 3Free University of Bozen-Bolzano Correspondence: leonardo.bertolazzi@unitn.it
@@ 1 Introduction
With the advent of increasingly capable large language models (LLMs), logical reasoning has become a central domain for evaluating and comparing these systems (Huang and Chang, 2023; Mondorf and Plank, 2024; Liu et al., 2025). However, language models’ ability to perform rigorous logical reasoning remains an open question. Even with extensive scientific, mathematical, and programming data, LLMs still fail to generalize beyond seen inference patterns (Clark et al., 2021; Saparov et al., 2023; Mirzadeh et al., 2025; Gulati et al., 2024; Huang et al., 2025; Shi et al., 2023; Yoran et al., 2024). These challenges relate to broader debates surrounding generalization versus memorization in LLMs (Balloccu et al., 2024; Singh et al., 2024).
Episode $\mathcal{T}$Knowledge Base ($\mathcal{KB}$)knowledge base: All x1 are x2, All x2 are x4, All x3 are x5, All x10 are x11, All x4 are x6, All x2 are x3, All x5 are x7, Some x5 are not x1, All x9 are x10, All x6 are x8, All x8 are x9, Some x11 are not x4 Study Examples ($S^{\text{supp}}$)<STUDY> hypothesis: All x8 are x11 premises: All x8 are x9, All x9 are x10, All x10 are x11; hypothesis: All x1 are x3 premises: All x1 are x2, All x2 are x3; … Query Hypothesis ($x^{\text{query}}$)<QUERY> hypothesis: All x3 are x7Query Premises ($y^{\text{query}}$)premises: All x3 are x5, All x5 are x7InputOutput
@@
Figure 1: Overview of a ML episode. Given a set of premises (the knowledge base, $\mathcal{KB}$), a set of task demonstrations (or Study Examples), and a Query Hypothesis $x^{\mathrm{query}}$ that is entailed from $\mathcal{KB}$, models must generate the minimal subset of premises, the Query Premises $y^{\mathrm{query}}$, from which $x^{\mathrm{query}}$ can be derived. During each ML episode, by being trained on the Study Examples, models learn to extract the abstract logical patterns. The examples show how we frame syllogistic inferences as a premise selection task. The dataset is built with pseudwords, where here we have variables for space reasons.
@@
Few-shot meta-learning approaches (Irie and Lake, 2024) have emerged as promising methods for inducing OOD generalization and rapid domain adaptation in LLMs. Specifically, this class of methods has proven effective in few-shot task generalization (Min et al., 2022; Chen et al., 2022), systematic generalization (Lake and Baroni, 2023), and mitigating catastrophic forgetting (Irie et al., 2025). In this work, we propose for the first time to apply few-shot meta-learning (ML) to enhance the ability of LLMs to learn logic.
@@
As illustrated in Figure 1, we evaluate the effectiveness of this approach using a logical reasoning task grounded in syllogistic logic (Smiley, 1973; Vargas Guzmán et al., 2024) and frame it as a premise selection task. This is a minimal but still interesting and non-trivial fragment of logic consisting of a few clear logical rules. As such it allows us to directly address our research question: Does ML enhance LLM’s ability to learn abstract logical patterns underlying reasoning and generalization? Each problem presents a knowledge base of atomic logical statements. Models are tasked with identifying the minimal subset of premises that logically entail a given test hypothesis. This premise selection task captures a core aspect of deductive reasoning: determining which known facts are necessary and sufficient to justify a conclusion. We apply ML to small LMs from the Qwen-2.5 family (Qwen Team, 2025), ranging from 1.5B to 7B parameters. Specifically, we assess the generalization capabilities induced by ML, such as systematically performing inferences over unseen sets of premises, as well as over more complex (longer) or simpler (shorter) sets of premises than those encountered during training. [^1] Our code and data are available at: https://github.com/leobertolazzi/meta-learning-logic.git.
Our main contributions are as follows: We introduce a new synthetic dataset based on syllogistic logic to directly study logical reasoning generalization in LLMs. We show that ML enables small LMs to better generalize in OOD syllogistic reasoning problems with particularly strong performance in smaller models and low-data regimes. We demonstrate that small LMs fine-tuned with ML can outperform state-of-the-art LLMs such as GPT-4o and o3-mini, on our syllogistic premise selection task.
@@ 2.1 Syllogistic Logic
Figure 2: Example inference. Edges labeled “All-are” denote universal affirmatives (e.g., All cats are felines). The solid red edge is a universal negative (No animals are plants). From these “atomic facts” we infer No cats are tulips (dashed red edge). Formally, this is expressed as $\{Aa-b,\;Ac-d,\;Ebd\}\vDash Eac$, the type 6 inference (Smiley, 1973). Here we use words to better explain the inference, the syntetic dataset models see consist of pseudowords.
@@
The language of syllogistic logic comprises a finite set of atomic terms $\{a,b,c,\ldots\}$ and four quantifier labels $A,E,I$, and $O$. Well-formed formulas consists of $Aab$ (“All $a$ are bb”), EabEab (“No $a$ are bb”), IabIab (“Some $a$ are bb”), and $Oab$ (“Some $a$ are not bb”). Finally, an A-chain, denoted as AabAa-b, represents the single formula $Aab$ or a sequence of formulas $Aac_{1}$, $Ac_{1}c_{2}$, $\dots$, $Ac_{n-1}c_{n}$, $Ac_{n}b$ for $n\geq 1$. A knowledge base ($\mathcal{KB}$) is defined as a finite set of formulas (premises).
@@
We aim for models to identify the minimal set of premises in a knowledge base to derive a given hypothesis. Formally, we are interested in inferences $\mathcal{F}\vDash F$ such that $\mathcal{F}^{\prime}\not\vDash F$ for any proper subset $\mathcal{F}^{\prime}\subsetneq\mathcal{F}$. For example, $\{Abc,Abd\}\vDash Icd$ is minimal, while $\{Aab,Abc,Abd\}\vDash Icd$ is not because $Aab$ is not needed to infer the conclusion.
There are seven types of minimal syllogistic inferences. To facilitate understanding, Figure 2 provides an intuitive representation of a type 6 inference. Further details about syllogistic logic can be found in Appendix A.
@@ 2.2 Meta-learning in Autoregressive Models
Meta-learning, or “learning to learn”, is a paradigm that aims to enable machine learning models to acquire transferable knowledge across multiple tasks, allowing rapid adaptation to new tasks with minimal data. Among the numerous existing meta-learning frameworks (Hospedales et al., 2022), our method is mainly inspired by Meta-learning Sequence Learners (MSL) (Irie and Lake, 2024).
@@
In standard supervised learning, data consists of a static dataset $\mathcal{D}_{\mathrm{train}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ where inputs $x_{i}$ are mapped to targets $y_{i}$ under a fixed distribution $p(x,y)$. By contrast, meta-learning organizes data into episodes $\mathcal{T}=(S^{\mathrm{supp}},S^{\mathrm{query}})$ drawn from $p(\mathcal{T})$, where $S^{\mathrm{supp}}=\{(x_{i},y_{i})\}_{i=1}^{K}$ is the support set containing task demonstrations, or study examples, and $S^{\mathrm{query}}=\{(x_{j},y_{j})\}_{j=1}^{M}$ is the query set for evaluation. We consider the simplest scenario where $|S^{\mathrm{query}}|=1$, containing a single example $(x^{\mathrm{query}},y^{\mathrm{query}})$. We adapt this episodic formulation to our task, as shown in Figure 1.
@@
The fundamental difference between the two paradigms appears in their optimization objectives. Standard supervised learning finds parameters $\theta^{*}$ that maximize the likelihood: $\theta^{*}=\mathrm{argmax}_{\theta}\sum_{(x,y)\in\mathcal{D}_{\mathrm{train}}}\log p_{\theta}(y\mid x)$. while meta-learning finds parameters $\theta^{*}$ that maximize the expected likelihood across tasks: $\theta^{*}=\mathrm{argmax}_{\theta}\quad\mathbb{E}_{\mathcal{T}}\left[\log p_{\theta}(y^{\mathrm{query}}\mid x^{\mathrm{query}},S^{\mathrm{supp}})\right]$.
For autoregressive models, the probability $p_{\theta}(y^{\mathrm{query}}\mid x^{\mathrm{query}},S^{\mathrm{supp}})$ is computed by conditioning on the support set $S^{\mathrm{supp}}$ as part of the input context, formatted as a sequence of input-output pairs preceding the query. This approach forces the model to develop the capabilities of recognizing and applying task patterns from the support examples to generate appropriate query outputs, namely, in our case, the minimal set of premises.
@@ 3.1 Data Generation
As illustrated in Figure 1, we structured each datapoint in the three parts to begin with the token “knowledge base:”, followed by the full sequence of premises, separated by commas. This is immediately followed by the special tag <QUERY>, and then the token “hypothesis:”, which introduces the target hypothesis. Next comes the token “premises:”, followed by the specific comma-separated premises that entail the hypothesis. To increase variability, we applied ten random pseudoword assignments and three random permutations of premise order for each $\mathcal{KB}$, resulting in multiple variants per datapoint.
@@
We generated enough $\mathcal{KB}$s to obtain 1000 training, 5 validation, and 100 test examples for each inference type and length combination in the range from 0 to 19.555Note that some inference types (e.g., type 3) span the full range of lengths from 0 to 19, while others span only a subrange (e.g., type 2 spans from 1 to 10). See all type-length combinations within the generated $\mathcal{KB}$s in Figure 7 in Appendix B.4. This range was chosen to allow experiments with generalization to both unseen shorter and longer inferences. Full dataset statistics, including the number of generated $\mathcal{KB}$s per split, are reported in Appendix B.4.
@@
3.2 Meta-learning for Logic
When applying meta-learning principles to the framework of syllogistic logic, we conceptualize the premises within a $\mathcal{KB}$ as atomic facts. The seven types of syllogism (as detailed in Table 4, in the Appendix) are treated as arguments, constructed using these atomic facts, and the model’s task is to extract the minimal set of facts within a $\mathcal{KB}$ to produce a valid argument that proves the query hypothesis.
The type of systematic generalization ML addresses involves applying the seven fixed syllogistic inferences to new, unseen sets of atomic facts. This is central to logical reasoning because logical rules are, by definition, formal: conclusions follow from premises based solely on the structure of the arguments, regardless of their specific content. Thus, successfully applying an inference type to a novel, unseen $\mathcal{KB}$ requires the model to recognize and instantiate the same formal structure with different premises. This generalization also includes variations in the number of atomic facts needed to instantiate an argument. Specifically, handling $A$-chains of varying lengths requires applying the learned inference patterns to longer or shorter instances of the same formal type.
@@ 3.2 Meta-learning for Logic
To induce meta-learning of inference types, ML uses a set of episodes where each episode $\mathcal{T}=(\mathcal{KB},S^{\mathrm{supp}},x^{\mathrm{query}},y^{\mathrm{query}})$. Here, $\mathcal{KB}$ is a knowledge base, $S^{\mathrm{supp}}$ is a set of study valid hypothesis-premises pairs, $x^{\mathrm{query}}$ is a valid query hypothesis, and $y^{\mathrm{query}}$ is the minimal set of premises, in the KB, entailing $x^{\mathrm{query}}$. Figure 1 shows a full ML episode using indexed variables in place of pseudowords for improved readability. Importantly, we consider study examples with inferences of the same type as the query. The number of study examples we set, i.e. valid hypothesis–premise pairs, is three. In their textual translation, we add the special tags <STUDY> to indicate the beginning of the sequence of study examples. During ML fine-tuning, models are trained to minimize the cross-entropy loss of the tokens in $y^{\mathrm{query}}$ given the input tokens from the context $(\mathcal{KB},S^{\mathrm{supp}},x^{\mathrm{query}})$. At inference time, ML models are provided as context $(\mathcal{KB},S^{\mathrm{supp}},x^{\mathrm{query}})$, and are tasked to generate $y^{\mathrm{query}}$.
@@
Similarly to Lake and Baroni (2023), we consider a baseline where models are fine-tuned only on single input-output pairs $(x^{\mathrm{query}},y^{\mathrm{query}})$ preceded by a $\mathcal{KB}$. The baseline is fine-tuned to minimize the cross-entropy loss of the tokens in $y^{\mathrm{query}}$ given the input tokens from the context $(\mathcal{KB},x^{\mathrm{query}})$. At inference time, we tested the baseline both with and without study examples from $S^{\mathrm{supp}}$ in the context. Across all experiments, we report the results for whichever setting performed best, with a direct comparison of both configurations provided in Appendix D.4. In the tables, we use subscript notation (e.g., Baseline+S or Baseline-S) to indicate whether the baseline was tested with or without study examples, respectively. To ensure a fair comparison between the meta-learning model and the baseline, we ensured that both models were fine-tuned on the exact same aggregate set of unique hypothesis-premises pairs. Specifically, the baseline was fine-tuned using a set $\mathcal{D}_{\text{baseline}}$ consisting of $(x^{\mathrm{query}},y^{\mathrm{query}})$ unique pairs. For the meta-learning approach, the corresponding set of all unique hypothesis-premises pairs encountered across all $N$ episodes $\mathcal{T}_{i}=(\mathcal{KB}_{i},S^{\mathrm{supp}}_{i},x^{\mathrm{query}}_{i},y^{\mathrm{query}}_{i})$ is given by $\mathcal{D}_{\text{meta}}=\bigcup_{i=1}^{N}(S^{\mathrm{supp}}_{i}\cup\{(x^{\mathrm{query}}_{i},y^{\mathrm{query}}_{i})\})$. We verified that $\mathcal{D}_{\text{baseline}}=\mathcal{D}_{\text{meta}}$. Moreover, since the meta-learning model processes more hypothesis-premises pairs within each training episode (due to $S^{\mathrm{supp}}_{i}$), we counterbalanced this by training the baseline model for a proportionally larger number of epochs. [^6] Further details on number of epochs for each approach are provided in Appendix C.2.
Practical considerations.
ML training episodes involve longer input sequences, which increase memory demands; to accommodate this, we applied quantization to the ML model so that all models can be trained and evaluated on a single 80GB A100 GPU. [^7] See Appendix C.2 for full details on weight precisions used in the experiments.